n!/2(n-1)!&n!/4!(n-4)!are in the ratio 2:1find n, x+yi=1+i whole square/1-i find x andy

Grade:12

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

n!/2(n-1)!&n!/4!(n-4)!are in the ratio 2:1

so, 12/(n-1)*(n-2)*(n-3) = 2

or, (n-1)*(n-2)*(n-3)= 6

so, n=4 (by hit n trial)

 

second part,  (1+i)2 =  1 - 1 +2i = 2i

so, 2i/1-i * 1+i/1+i = 2i -2/2 = i -1 ( rationalize LHS)

now, x +iy = i -1

so, x = -1, y =1 (equate coefficient )

 

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..